The velocity of a falling object is a linear function of the time t it has been falling. If v = 20 when t = 0 and
v = 53 when t = 11, express v as a function of t.



Answer by ichudov(507)

We know that velocity is V = a+bt, where a is a constant and t is the variable.
When t = 0, a+b*0 = 20
When t = 11, a+b*11 = 53
This is a simple linear system, which you can solve using a linear system solver that will show work. The solution is pretty obvious, the first equation gives you a=20, and the second equation b*11 = 33, gives you b = 11.